# LCI: a Lightweight Communication Interface for Efficient Asynchronous Multithreaded Communication

### v1

@@ Abstract.
The evolution of architectures, programming models, and algorithms is driving communication towards greater asynchrony and concurrency, usually in multithreaded environments. We present LCI, a communication library designed for efficient asynchronous multithreaded communication. LCI provides a concise interface that supports common point-to-point primitives and diverse completion mechanisms, along with flexible controls for incrementally fine-tuning communication resources and runtime behavior. It features a threading-efficient runtime built on atomic data structures, fine-grained non-blocking locks, and low-level network insights. We evaluate LCI on both Inifiniband and Slingshot-11 clusters with microbenchmarks and two application-level benchmarks. Experiment results show that LCI significantly outperforms existing communication libraries in various multithreaded scenarios, achieving performance that exceeds the traditional multi-process execution mode and unlocking new possibilities for emerging programming models and applications.
@@
copyright: none
@@ 1. Introduction
These characteristics fall out of the original focus of MPI, the de facto standard HPC communication library designed over 30 years ago. Since then, new communication libraries and MPI features have been introduced to tackle the asynchrony. Multiple research efforts, mainly by the MPI community, have been taken to improve multithreaded communication support. However, they still fall short of the needs of applications due to limited flexibility and constrained design space.
  • Limited Flexibility: Each communication library only offers a limited selection of communication mechanisms. However, modern programming systems and/or applications can need combinations of many communication mechanisms. Clients often must implement their communication mechanisms on top of the existing library interface. This requires a significant effort and is not optimal when the library does not expose low-level functionality.
@@
To address these issues, we present the Lightweight Communication Interface (LCI), a communication library designed from scratch with asynchronous multithreaded communication in mind. It provides a unified interface that supports flexible combinations of all common point-to-point communication primitives, including send-receive, active messages, and RMA put/get (with/without notification), and various built-in mechanisms to synchronize with pending communications, including synchronizers, completion queues, function handlers, and completion graphs. In addition, the interface offers both a simple starting point for users to program and a wide range of options for them to incrementally fine-tune the communication resources and runtime behaviors, minimizing potential interference between communication and computation. Finally, it is supported by a lightweight and efficient runtime optimized for threading efficiency and massive parallelism. The runtime is built with a deep understanding of low-level network activities and employs optimizations such as atomic-based data structures, thread-local storage, and fine-grained nonblocking locks.
@@
The rest of the paper is organized as follows: Section 2 discusses related works. Section 3 introduces LCI’s communication interface and shows how it can seamlessly support dynamic programming systems. Section 4 presents the key designs in LCI runtime. Section 5 analyzes the evaluation results. Section 6 concludes the paper and discusses future works.
@@ 2.1. Asynchronous Communication
New MPI features have been proposed since then to improve support for asynchronous communication. RMA operations have been included in MPI since MPI-2, though its ”window” abstraction still operates in a partially collective style. The MPI continuation proposal (Schuchart et al., 9 01) recently introduced a way for clients to attach callbacks to pending MPI operations, aiming for more efficient polling in the case of heavy communication overlapping.
A range of communication libraries has also been proposed. GASNet(Bonachea and Jeong, 2002) and the later GASNet-EX (Bonachea and Hargrove, 2018) focus on one-sided active messages and RMA primitives. They are intended to be used by runtime developers or as a compilation target, so their interfaces are generally more complicated than MPI. At a higher level, PGAS libraries/languages, such as UPC (El-Ghazawi and Smith, 2006), UPC++ (Bachan et al., 2019), and OpenSHMEM(Chapman et al., 2010), rely on RMA operations to maintain a global address abstraction. Recently, new communication libraries have been proposed. YGM (Steil et al., 2023) features a batch-processing active messages interface and utilizes aggregation for better throughput. UNR (Feng et al., 2024) emphasizes notifiable RMA operations, optimizing them for multi-NIC aggregation and ease of use.
UCX(Shamis et al., 2015) and Libfabric((2024), OFIWG) provide low-level abstractions that are portable across multiple interconnects. They offer more flexible interfaces but at a much lower level. They also require manual bootstrap. Their primary usage is to support communication libraries rather than high-level programming systems/applications.
While these libraries have made significant progress in supporting asynchronous communication, they often provide a limited selection of features that cannot fully fulfill the communication needs of complicated runtime systems/applications. LCI improves upon these libraries by providing a more comprehensive and flexible interface that allows for a broader range of communication patterns and optimizations. It also has an additional performance focus on multithreaded communication. (Yan and Snir, 2025) presents an overview and some considerations of an earlier version of the LCI interface in a workshop paper.
@@ 2.2. Multithreaded Communication
All major communication libraries can be configured to be thread-safe, but the resulting performance is often suboptimal. The work on MPI and GASNet started when processors had a single core, so multithreading was not a concern. Some aspects of the interface design proved problematic when multithreading was retrofitted. Furthermore, as early applications were single-threaded, MPI implementers focused on single-threaded performance. Consequently, users kept communication single-threaded (one process per core or one communication thread per process model), reinforcing the emphasis on single-threaded performance.
@@
Our work builds upon the valuable insights of existing works and advances them through completely redesigning the communication interface and runtime, free from backward compatibility concerns. By adopting appropriate interface options and semantics, decomposing the runtime into multiple independent resources, and applying various optimization techniques, we present a communication library that, for the first time known to us, achieves multithreaded performance surpassing multi-process performance at the microbenchmark level.
@@ 3. LCI Interface
The LCI interface is designed to be intuitive, flexible, and explicit, allowing LCI to be seamlessly integrated into complicated runtimes with diverse communication needs. We first present the Objectifed Flexible Functions (OFF) idiom that allows users to specify optional arguments in any order as a C++ function call. All LCI functions have a variant adopting this idiom. We then walk through the core LCI interface by building an LCI backend for a simple Remote Procedure Call (RPC) library. Finally, we discuss other important details of the LCI interface.
3.1. Objectified Flexible Function
The LCI interface is designed to be flexible and customizable, allowing users to express their communication needs in a straightforward and efficient manner. As a result, some LCI operations have many optional arguments. The C++ optional argument semantic is not flexible enough to handle them, as it only allows users to specify the optional arguments in the order they are defined with no gaps.
We propose a new C++ idiom called Objectified Flexible Function (OFF) to overcome this restriction. It allows users to specify the optional arguments in any order, similar to the named optional arguments in Python functions. Listing 1 shows what it is like to invoke an OFF operation using the post_send operation in LCI as an example. Line 1 invokes the post_send operation in its standard form with only the positional arguments. Lines 2-3 invoke the OFF variant of the same operation. Line 2 associates the send with a specific device, and Line 3 further specifies the rank_only matching policy. The OFF variant in LCI is always suffixed with _x.
1auto ret = post_send(rank, buf, size, tag, comp);
2auto ret = post_send_x(rank, buf, size, tag, comp).device(device)();
3auto ret = post_send_x(rank, buf, size, tag, comp).matching_policy(matching_policy_t::rank_only).device(device)();
Listing 1: Objectified Flexible Function Example.
The OFF idiom allows LCI to maintain the API’s conciseness while providing as much flexibility as possible. The user can start from the simplest form and incrementally refine the communication behavior in any direction they need.
Under the hood, an OFF is implemented as a functor with a constructor that takes the positional arguments and a set of setter methods for the optional arguments. We use a Python script to generate the OFF definition based on a DSL input.
@@ 3.2.1. The iRPCLib Example
Remote Procedure Calls (RPCs) are a popular programming paradigm that allows a client to invoke arbitrary functions on a server. The main difference between RPC and active message is that the active message handler is executed inside the low-level communication progress engine and thus is supposed to be short with restricted functionalities (e.g., cannot invoke another communication). In contrast, RPC handlers usually have no restrictions. RPCs are used extensively in high-level programming models (Bachan et al., 2019; Kaiser et al., 2023; Moritz et al., 2018). This section illustrates the LCI interface by building an LCI backend for an imaginary RPC library (iRPCLib).
@@
6__thread lci::device_t device;
@@
18 shandler = lci::alloc_handler(upper_layer::send_cb);
@@
41 if (status.error.is_done())
@@
44 assert(status.error.is_posted());
@@
52 if (status.error.is_done()) {
53 lci::buffer_t buf = status.get_buffer();
@@
55 .rank = status.rank,
56 .tag = status.tag,
57 .buf = buf.base,
58 .size = buf.size,
@@
64 assert(status.error.is_retry());
@@
68
69bool do_background_work() {
70 return lci::progress_x().device(device)();
71}
@@
Listing 2 shows the example implementation of the iRPCLib LCI backend. We assume iRPCLib has two layers, the upper layer and the backend layer. The upper layer is responsible for registering the user-provided RPC handlers into indices and serializing and deserializing the RPC arguments into consecutive memory buffers (not shown here). The backend layer is responsible for sending the RPC handler index (tag) and serialized arguments (pointed by buf) to the target rank (send_msg in Line 37) and deliver the incoming messages to the upper layer (poll_msg in Line 50). For simplicity, we assume iRPCLib just wants the backend layer to free the message buffer once the send completes locally (send_cb in Line 9). We further assume iRPCLib is multithreaded. The main thread will call global_init (Line 14) and global_fina (Line 23) and all threads will call thread_init (Line 29) and thread_fina (Line 33) during the initialization and finalization phases. All threads can produce and consume communication (a.k.a. calling send_msg and poll_msg). In addition, all threads will periodically call do_background_work (Line 69) to make progress on the pending communication. The backend abstraction described here is a simplified version of the HPX parcelport abstraction (Yan et al., 2025) and the Charm++ Converse Machine Interface (Kalé et al., 1996).
@@ 3.2.2. Runtime Lifecycle
LCI does not have global initialization or finalization functions. Instead, it provides functions to (de)allocate a runtime object. The runtime object wraps default configurations and communication resources for LCI to operate. Most LCI operations accept runtime as an optional argument. In Listing 2, iRPCLib just uses the global default runtime (g_runtime) for simplicity (Lines 15, 26). Once at least one runtime is active, the user can query the rank of the current process (Line 16) and the total number of ranks (Line 17).
An LCI client typically allocates only one runtime object. However, multiple runtime objects can exist due to library composition. In these cases, the runtime abstraction enables different libraries to use different configurations and resources without interfering with each other.
@@ 3.2.3. Resource
Communications operate on resources. LCI allows users to allocate resources explicitly and associate them with communications. Resources can have a list of attributes. Users can explicitly set them during resource allocation and query them afterward. In Listing 2, iRPCLib uses one device per thread to improve threading efficiency (Line 6). A device encapsulating a complete set of low-level network resources and LCI ensures threads operating on different devices will not interfere with each other. In addition, iRPCLib uses a shared completion handler (shandler in Line 2) for source completion and a shared completion queue (rcq in Line 3) for target completion. Line 20 further registers the completion queue into a remote completion handle (rcomp) for other processes to post active messages to. (See Section 3.2.5 for more detail.)
Other important LCI resources (not shown here) include (a) matching engines matching send and receive; (b) packet pools (de)allocating fixed-sized pre-registered internal buffers (packets); and (c) backlog queue storing temporarily postponed communication requests. A communication operation is free to associate with any combination of these resources. For example, if the iRPCLib also uses send-receive, all threads can use a shared matching engine while using per-thread devices. In this way, it could achieve great threading efficiency while maintaining a global matching domain. Section 4.1 talks about resources in more detail.
@@ 3.2.4. Communication Posting
Line 38 uses the LCI active message operation to send the message along with a tag to the target rank using the thread-local device. LCI supports all commonly used point-to-point communication paradigms, including send/receive, active message, and RMA put/get. It supports them in a unified manner to reduce the API’s complexity and allow users to easily switch between different communication models.
@@
The parameters needed to specify a communication are mostly the same across all point-to-point communication paradigms. Different communication paradigms are just different choices of where to specify these parameters. For example, send-recv specifies only the local parameters on each side, while RMA put/get specifies all parameters on only one side.
Therefore, LCI offers a generic communication posting operation, post_comm. This operation takes the target rank, the local buffer, the message size, and the local completion object as positional arguments. It takes a wide range of optional arguments, among which the most important ones include the direction, the remote buffer, and the remote completion object. Table 1 shows how combining the three optional arguments can specify the common point-to-point communication paradigms.

### v2

@@ Abstract.
The evolution of architectures, programming models, and algorithms is driving communication towards greater asynchrony and concurrency, usually in multithreaded environments. We present LCI, a communication library designed for efficient asynchronous multithreaded communication. LCI provides a concise interface that supports common point-to-point primitives and diverse completion mechanisms, along with flexible controls for incrementally fine-tuning communication resources and runtime behavior. It features a threading-efficient runtime built on atomic data structures, fine-grained non-blocking locks, and low-level network insights. We evaluate LCI on both Infiniband and Slingshot-11 clusters with microbenchmarks and two application-level benchmarks. Experimental results show that LCI significantly outperforms existing communication libraries in various multithreaded scenarios, achieving performance that exceeds the traditional multi-process execution mode and unlocking new possibilities for emerging programming models and applications. LCI is open-source and available at https://github.com/uiuc-hpc/lci.
@@
journalyear: 2025copyright: ccconference: The International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; StLouis, MO, USAbooktitle: The International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3712285.3759881isbn: 979-8-4007-1466-5/2025/11ccs: Networks Programming interfacesccs: Computing methodologies Parallel programming languages
@@ 1. Introduction
These characteristics fall out of the original focus of MPI, the de facto standard HPC communication library designed over 30 years ago. Since then, new communication libraries and MPI features have been introduced to enhance support for asynchrony. Multiple research efforts, mainly by the MPI community, have also attempted to address the challenges of multithreaded communication. However, they still fall short of the needs of applications due to limited flexibility and constrained design space.
  • Limited Flexibility: Each communication library only offers a limited selection of communication mechanisms. However, modern programming systems and/or applications can need combinations of many communication mechanisms. Clients often must implement their communication mechanisms on top of the existing library interface. This requires a significant effort and is not optimal when the library does not expose low-level network functionality.
@@
To address these issues, we present the Lightweight Communication Interface (LCI), a communication library designed from scratch with asynchronous multithreaded communication in mind. It provides a unified interface that supports flexible combinations of all common point-to-point communication primitives, including send-receive, active messages, and RMA put/get (with/without notification), and various built-in mechanisms to synchronize with pending communications, including counters, synchronizers, completion queues, function handlers, and completion graphs. In addition, the interface offers both a simple starting point for programming and a wide range of options to incrementally fine-tune the communication resources and runtime behaviors, minimizing interference between communication and computation. Finally, it is backed by a lightweight and efficient runtime optimized for threading efficiency and massive parallelism. The runtime is built with a deep understanding of low-level network activities and employs optimizations such as atomic-based data structures, thread-local storage, and fine-grained nonblocking locks.
@@
The rest of the paper is organized as follows: Section 2 discusses related works. Section 3 provides an overview of LCI. Section 4 introduces LCI’s communication interface and shows how it can seamlessly support dynamic runtime systems. Section 5 presents the key designs in the LCI runtime. Section 6 analyzes the evaluation results. Section 7 concludes the paper and discusses future work.
@@ 2.1. Asynchronous Communication
Since then, new MPI features have been proposed to improve support for asynchronous communication. RMA operations have been included in MPI since MPI-2, though its ”window” abstraction still operates in a partially collective style. The MPI continuation proposal (Schuchart et al., 9 01) recently introduced a way for clients to attach callbacks to pending MPI operations, aiming for more efficient polling in the case of heavy communication overlapping.
Other communication libraries have also been proposed. GASNet (Bonachea and Jeong, 2002) and the later GASNet-EX (Bonachea and Hargrove, 2018) focus on one-sided active messages and RMA primitives. They are intended to be used by runtime developers or as a compilation target, so their interfaces are more complicated than MPI. At a higher level, PGAS libraries and languages, such as UPC (El-Ghazawi and Smith, 2006), UPC++ (Bachan et al., 2019), and OpenSHMEM(Chapman et al., 2010), rely on RMA operations to implement a global address abstraction. Recently, new communication libraries have been proposed. YGM (Steil et al., 2023) features a batch-processing active messages interface and utilizes aggregation for better throughput. UNR (Feng et al., 2024) emphasizes notifiable RMA operations, optimizing them for multi-NIC aggregation and ease of use.
UCX (Shamis et al., 2015) and Libfabric ((2024), OFIWG) provide portable low-level abstractions across multiple interconnects. They offer more flexible interfaces, but at a much lower level. They also require manual bootstrapping. Their primary usage is to support communication libraries rather than high-level programming systems or applications.
While these libraries have made significant progress in supporting asynchronous communication, they often provide a limited selection of features that cannot fully fulfill the communication needs of complicated runtime systems or applications. LCI improves upon these libraries by providing a more comprehensive and flexible interface that allows for a broader range of communication patterns and optimizations. It also has an additional performance focus on multithreaded communication.
Earlier work on LCI includes its integration into PaRSEC (Mor et al., 2023), resulting in a significant improvement in the performance of the HiCMA sparse Cholesky solver, and its integration into HPX releases (Yan et al., 2023, 2025a). (Strack et al., 2024) used such a release to implement a 2D FFT mini-app with HPX that outperforms the MPI counterpart and the reference FFTW implementation by 5x. (Yan and Snir, 2025) presents an initial overview and some considerations of the LCI interface in a workshop paper. This paper is the first full publication dedicated to LCI’s interface and runtime design.
@@ 2.2. Multithreaded Communication
All major communication libraries can be configured to be thread-safe (e.g., MPI_THREAD_MULTIPLE), but the resulting performance is often suboptimal. The work on MPI and GASNet started when processors had a single core, so multithreading was not a concern. Some aspects of the interface design proved problematic when multithreading was retrofitted. Furthermore, as early applications were single-threaded, MPI implementers focused on single-threaded performance. Consequently, users kept communication single-threaded (one process per core or one communication thread per process model), reinforcing the emphasis on single-threaded performance.
@@
Our work builds upon the valuable insights of existing works and advances them through completely redesigning the communication interface and runtime, free from backward compatibility concerns. By adopting appropriate interface options and semantics, decomposing the runtime into multiple independent components (referred to in LCI as resources), and applying various optimization techniques, we present a communication library that, to the best of our knowledge, is the first to achieve multithreaded performance surpassing multi-process performance in pure communication microbenchmarks.
3. LCI Overview
Figure 1. LCI Overview.
Figure 1 shows the high-level architecture of LCI. LCI is a high-level communication library for applications and runtimes with dynamic communication patterns. The LCI interface provides a set of operations for communication posting, completion checking, progressing, and resource Management. The LCI runtime executes these communication operations and is built on top of multiple internal components (resources), including network devices, matching engines, packet pools, completion objects, and backlog queues. The runtime relies on optimized multi-producer-multi-consumer (MPMC) data structures for efficient threading. LCI supports multiple bootstrap mechanisms, including PMI1, PMI2, PMIx, MPI, and Linux flock (file lock), allowing it to operate in a wide range of environments. For the network backends, it supports libibverbs for Infiniband and RoCE and libfabric for Slingshot-11 and other interconnects.
@@ 4. LCI Interface
The LCI interface is designed to be intuitive, flexible, and explicit, allowing LCI to be seamlessly integrated into complicated runtimes with diverse communication needs. We first introduce the Named Parameter Idiom, which allows users to specify optional arguments in any order in a C++ function call. All LCI functions have variants using this idiom. We then walk through the core LCI interface by building an LCI backend for a simple Remote Procedure Call (RPC) library. Finally, we discuss other important details of the LCI interface.
4.1. Named Parameter Idiom
The LCI interface is designed to be flexible and customizable. As a result, some LCI operations have many optional arguments. However, C++’s default optional argument semantics are not flexible enough to handle them, as they only allow users to specify the optional arguments in the order they are defined, with no gaps.
LCI adopts the C++ Named Parameter Idiom (C++, [n. d.]) to overcome this limitation. It allows users to specify the optional arguments in any order with their names. Listing 1 illustrates this idiom using the post_send operation in LCI as an example. Line 1 invokes the post_send operation in its standard form with only the positional arguments. Lines 2-3 invoke the named-parameter variant of the same operation. Line 2 associates the send with a specific device, and Line 3 further specifies a non-default matching engine.
The named-parameter variants in LCI are suffixed with _x. Unlike the original idiom, we use the () operator, instead of a special constructor of the returned object, to invoke the actual operation. This approach retains the ability to return scalar values directly from the operations.
1status = post_send(rank, buf, size, tag, comp);
2status = post_send_x(rank, buf, size, tag, comp).device(device)();
3status = post_send_x(rank, buf, size, tag, comp).matching_engine(matching_engine).device(device)();
Listing 1: Named Parameter Idiom Example.
The Named Parameter idiom allows LCI to maintain the API’s conciseness while providing maximal flexibility. The user can start with the simplest form and incrementally refine the communication behavior in any direction they need.
@@ 4.2.1. The iRPCLib Example
Remote Procedure Calls (RPCs) are a popular programming paradigm allowing clients to invoke arbitrary functions on a server. The main difference between RPCs and active messages is that the active message handler is executed inside the low-level network context and thus is supposed to be short with restricted functionalities (e.g., cannot invoke another communication). In contrast, RPC handlers usually have no restrictions. RPCs are used extensively in high-level programming models (Bachan et al., 2019; Kaiser et al., 2023; Moritz et al., 2018). This section illustrates the LCI interface by building an LCI backend for an imaginary RPC library (iRPCLib).
@@
6thread_local lci::device_t device;
@@
18 shandler = lci::alloc_handler(send_cb);
@@
41 else if (status.is_done())
@@
44 assert(status.is_posted());
@@
52 if (status.is_done()) {
@@
54 .rank = status.get_rank(),
55 .tag = status.get_tag(),
56 .buf = status.get_buffer(),
57 .size = status.get_size(),
@@
63 assert(status.is_retry());
@@
68void do_background_work() {
69 lci::progress_x().device(device)();
70}
@@
Listing 2 shows the example implementation of the iRPCLib LCI backend. We assume iRPCLib has two layers, the upper layer and the backend layer. The upper layer is responsible for registering the user-provided RPC handlers into indices and serializing and deserializing the RPC arguments into consecutive memory buffers (not shown here). The backend layer is responsible for sending the RPC handler index (tag) and serialized arguments (pointed to by buf) to the target rank (send_msg in Line 37) and delivering the incoming messages to the upper layer (poll_msg in Line 50). For simplicity, we assume iRPCLib just wants the backend layer to free the message buffer once the send completes locally (send_cb in Line 9). We further assume iRPCLib is multithreaded. The main thread will call global_init (Line 14) and global_fina (Line 23) and all threads will call thread_init (Line 29) and thread_fina (Line 33) during the initialization and finalization phases. All threads can produce and consume communication (a.k.a. calling send_msg and poll_msg). In addition, all threads will periodically call
do_background_work (Line 68) to make progress on the pending communication. The backend abstraction described here is conceptually a simplified version of the HPX parcelport abstraction (Yan et al., 2025a) and the Charm++ Converse Machine Interface (Kale et al., 1996).
@@ 4.2.2. Runtime Lifecycle
LCI does not have global initialization or finalization functions. Instead, it provides functions to (de)allocate a runtime object. The runtime object wraps default configurations and communication resources for LCI to operate. Most LCI operations accept runtime as an optional argument. In Listing 2, iRPCLib just uses the global default runtime (g_runtime) for simplicity (Lines 15, 26). Once at least one runtime is active, the user can query the rank of the current process (Line 16) and the total number of processes (Line 17).
An LCI client typically allocates only one runtime object. However, multiple runtime objects can exist due to library composition. In these cases, the runtime abstraction enables different libraries to use distinct configurations and resources without interfering with one another.
@@ 4.2.3. Resource
Communications operate on resources. LCI allows users to allocate resources explicitly and associate them with communications. Resources can have a list of attributes. Users can explicitly set them during resource allocation and query them afterward. In Listing 2, iRPCLib uses one device per thread to improve threading efficiency (Line 6). A device encapsulating a complete set of low-level network resources and LCI ensures threads operating on different devices will not interfere with each other. In addition, iRPCLib uses a shared completion handler (shandler in Line 2) for source completion and a shared completion queue (rcq in Line 3) for target completion. Line 20 further registers the completion queue into a remote completion handle (rcomp) for other processes to post active messages to. (See Section 4.2.6 for more details.)
Other important LCI resources (not shown here) include (a) matching engines matching send and receive; (b) packet pools (de)allocating fixed-sized pre-registered internal buffers (packets); and (c) backlog queue storing temporarily postponed communication requests. A communication operation is free to associate with any combination of these resources. For example, if the iRPCLib also uses send-receive, all threads can use a shared matching engine while using per-thread devices. This way, it could achieve great threading efficiency while maintaining a global matching domain. Section 5.1 discusses resources in more detail.
@@ 4.2.4. Communication Posting
Line 38 uses the LCI active message operation to send the message along with a tag to the target rank using the thread-local device. LCI supports all commonly used point-to-point communication paradigms, including send/receive, active message, and RMA put/get. It supports them in a unified manner to reduce the API’s complexity and allow users to switch between different communication models easily.
@@
The parameters needed to specify a communication are mostly the same across all point-to-point communication paradigms. The communication paradigm dictates which side specifies the parameters. For example, send-recv specifies only the local parameters on each side, while RMA put/get specifies all parameters only on the source/target side.
Therefore, LCI offers a generic communication posting operation, post_comm. This operation takes the target rank, the local buffer, the message size, and the local completion object as positional arguments. It takes a wide range of optional arguments, among which the most important ones include the direction, the remote buffer, and the remote completion object. Table 1 shows how combining the three optional arguments can instantiate common point-to-point communication paradigms.